# Charge-ordered states in twisted MoTe2

### 05 Mar 2026 v1

@@ I.3 Effective potential
$\Pi_{+}=i\frac{\sqrt{2}\hbar}{l_{B}}\,a^{\dagger},\qquad\Pi_{-}=-i\frac{\sqrt{2}\hbar}{l_{B}}\,a.$ (S23)
@@
$\langle 1X^{\prime}|e^{i\bm{G}\cdot\bm{r}}|0X\rangle=i\frac{G_{-}l_{B}}{\sqrt{2}}\,\langle 0X^{\prime}|e^{i\bm{G}\cdot\bm{r}}|0X\rangle,$ (S24)
@@ II Landau-level operators
The following relations are useful and consistent with electrons in a magnetic field $\bm{B}=(0,0,-|B|)$ (equivalently, holes in $\bm{B}=(0,0,|B|)$) [60]:
$\Pi_{x}$ $=p_{x}+eA_{x},\quad\Pi_{y}=p_{y}+eA_{y},\quad[\Pi_{i},\Pi_{j}]=+i\hbar^{2}l_{B}^{-2}\epsilon_{ij},\quad l_{B}=\sqrt{\hbar/eB},$ (S28)
$a$ $=i\frac{l_{B}}{\hbar\sqrt{2}}\left(\Pi_{x}+i\Pi_{y}\right),\quad a^{\dagger}=-i\frac{l_{B}}{\hbar\sqrt{2}}\left(\Pi_{x}-i\Pi_{y}\right),\quad[a,a^{\dagger}]=1,[a^{\dagger}a,a^{\dagger}]=a^{\dagger},$ (S29)
$\Pi_{x}$ $=i\frac{\hbar}{l_{B}\sqrt{2}}(a-a^{\dagger}),\quad\Pi_{y}=\frac{\hbar}{l_{B}\sqrt{2}}(a^{\dagger}+a),$ (S30)
$H_{0}$ $=\hbar\omega_{c}\left(a^{\dagger}a+\frac{1}{2}\right),\quad\omega_{c}=eB/m,$ (S31)
$\bm{R}$ $=\bm{r}-\frac{l_{B}^{2}}{\hbar}\hat{z}\times\bm{\Pi},\quad[R_{i},R_{j}]=-il_{B}^{2}\epsilon_{ij},\quad[R_{i},\Pi_{j}]=0,\quad[R_{i},H_{0}]=0,$ (S32)
$b$ $=\frac{1}{\sqrt{2}l_{B}}(R_{x}+iR_{y}),\quad b^{\dagger}=\frac{1}{\sqrt{2}l_{B}}(R_{x}-iR_{y}),\quad[b,b^{\dagger}]=1,\quad[b,a]=[b^{\dagger},a]=0,$ (S33)
$\bm{K}$ $=\bm{\Pi}+\hbar l_{B}^{-2}(\hat{z}\times\bm{r})=+\hbar l_{B}^{-2}(\hat{z}\times\bm{R}),\quad[K_{i},K_{j}]=-i\epsilon_{ij}\hbar^{2}l_{B}^{-2},$ (S34)
@@ III Commensurability of CDW and moiré patterns in a magnetic field
Let $n_{\phi}^{M}$ and $n_{\phi}^{\mathrm{CDW}}$ denote the magnetic flux per moiré and CDW unit cell, respectively. The topological non-trivial winding of the texture leads to $n_{\phi}^{M}=1$. We focus on commensurate periodicity that break translation symmetry while preserving $C_{6}$ rotational symmetry, implying a larger real-space CDW unit-cell area and $n_{\phi}^{\mathrm{CDW}}>n_{\phi}^{M}$. Assuming rational flux per CDW unit cell,
@@ IV Particle–hole transformation
$H_{\text{ad}}\rightarrow H_{\text{ad}}$ $=\frac{(\hat{\bm{p}}+e\bm{\mathcal{A}})^{2}}{2m}-V^{*}(\bm{r})$ (S45)
$=\frac{(\hat{\bm{p}}+e\bm{\mathcal{A}}_{0})^{2}}{2m}+\frac{1}{2m}\{(\hat{\bm{p}}+e\bm{\mathcal{A}}_{0}),\delta\bm{\mathcal{A}}(\bm{r})\}+\frac{(\delta\bm{\mathcal{A}}(\bm{r}))^{2}}{2m}-U_{\text{ad}}(\bm{r}).$ (S46)
@@ V Hartree–Fock matrix elements
We evaluate the exchange term analytically and implement the resulting closed-form expression in numerics [31] (Note, numerical implementations can be also be found [33]). Total, single-particle, and interaction energies are
$E_{\mathrm{sp}}\equiv\mathrm{tr}(H^{0}\rho),\qquad E_{\mathrm{tot}}=\mathrm{tr}(H^{0}\rho)+\frac{1}{2}\,\mathrm{tr}(\Sigma^{\rm HF}\rho),\qquad E_{\mathrm{int}}=E_{\mathrm{tot}}-E_{\mathrm{sp}}.$ (S57)
@@ V.1 Density matrix and relations
where we used $\langle n^{\prime}X^{\prime}|e^{i\bm{q}\cdot\bm{r}}|nX\rangle=\delta_{X^{\prime},\,X+\ell_{B}^{2}q_{y}}\,e^{iq_{x}(X+X^{\prime})/2}\,F_{n^{\prime}n}(\bm{q})$ (see above) and the the guiding-center density matrices
@@
[…]
@@ V.2 Sum rule for the lowest Landau level
the usual LLL sum rule [31, 58, 57]. With $L$ mixing, $\sum_{\bm{G}}|\Delta_{00}(\bm{G})|^{2}<\nu$.
@@ VI Lowest-Landau-level weight
To quantify how much a self-consistent solution resides in the LLL, define the lowest landau level weight $w_{\mathrm{LLL}}$:
@@ VII.1 Plane-wave matrix elements
$\langle k_{x}^{\prime}k_{y}^{\prime};n^{\prime}i^{\prime}|e^{i\bm{G}\cdot\bm{r}}|k_{x}k_{y};ni\rangle$ $=\delta_{k_{y},k_{y}^{\prime}}\,\delta_{k_{x},k_{x}^{\prime}}\,F_{n^{\prime}n}(\bm{G})\,e^{iQ_{2}\,b\,k_{x}\ell_{B}^{2}}\,e^{i\ell_{B}^{2}G_{x}(k_{y}+G_{y}/2)}\,e^{iG_{x}\ell_{B}^{2}Q_{2}i}\,\delta_{\,i^{\prime},\,(i+b)\bmod j}.$ (S71)
@@
$\langle k_{x}^{\prime}k_{y}^{\prime};n^{\prime}i^{\prime}|e^{i\bm{G}\cdot\bm{r}}|k_{x}k_{y};ni\rangle$ $=\delta_{k_{y},k_{y}^{\prime}}\,\delta_{k_{x},k_{x}^{\prime}}\,\delta_{\,i^{\prime},\,(i+b)\bmod j}\,F_{n^{\prime}n}(\bm{G})\,e^{2\pi i\,b\,\tilde{k}_{x}/j}\,e^{2\pi i\,\tilde{G}_{x}(\tilde{k}_{y}+b/2)/n_{\phi}}\,e^{2\pi i\,\tilde{G}_{x}i/n_{\phi}}.$ (S72)
@@
$N_{y}=\frac{L}{Q_{2}\,j\,\ell_{B}^{2}}.$ (S73)
We use $\mathcal{N}$ solely for state normalization, and $N_{y}$ for counting magnetic unit cells.
@@ VII.3 Chern number
$|u_{\bm{k}};n,i\rangle=e^{-i\bm{k}\cdot\hat{\bm{r}}}|\bm{k};n,i\rangle\qquad\Longrightarrow\qquad\langle u_{\bm{k}};n,i|u_{\bm{k}^{\prime}};n^{\prime},i^{\prime}\rangle=\langle\bm{k};n,i|e^{i(\bm{k}-\bm{k}^{\prime})\cdot\bm{r}}|k^{\prime};n^{\prime},i^{\prime}\rangle.$ (S74)
@@
$\langle\bm{k};n,i|e^{i(\bm{k}-\bm{k}^{\prime})\cdot\bm{r}}|k^{\prime};n^{\prime},i^{\prime}\rangle$ $=\delta_{i,i^{\prime}}F_{n^{\prime}n}(\bm{k}-\bm{k}^{\prime})e^{-i\pi(\tilde{k}_{x}+\tilde{k}_{x}^{\prime})(\tilde{k}_{y}-\tilde{k}_{y}^{\prime})/j},\quad k_{x}=\tilde{k}_{x}Q_{0}p/(qj),\quad k_{y}=\tilde{k}_{y}Q_{2}.$ (S75)
@@
$|\bm{k},\alpha\rangle=\sum_{n,i}C^{\alpha}_{n,i}|\bm{k};n,i\rangle,$ $|u_{\bm{k}\alpha}\rangle=\sum_{n,i}C^{\alpha}_{n,i}e^{i\bm{k}\cdot\hat{\bm{r}}}|\bm{k};n,i\rangle,$ $\Longrightarrow$ $\langle u_{\bm{k}\alpha}|u_{\bm{k}^{\prime}\alpha^{\prime}}\rangle=(\bm{C}^{\alpha}_{\bm{k}})^{\dagger}\mathrm{PW}(\bm{k},\bm{k}^{\prime})\bm{C}^{\alpha^{\prime}}_{\bm{k}^{\prime}},$ (S76)
@@
VIII Dielectric constant dependent phase transition at $\nu_{h}=3/4$
Here we discuss the dielectric-constant dependence of the ground-state energy at $\nu_{h}=3/4$. We find a pronounced $\varepsilon$-dependence of the energetically preferred lattice pinning configuration. For stronger Coulomb interactions (smaller $\varepsilon$), the $C_{6}$-symmetry-broken state in which the electron is pinned to the MX or XM site has lower energy for all twist angles in $2.5^{\circ},4.5^{\circ}]$, whereas for weaker interactions (larger $\varepsilon$) the transition shifts towards smaller twist angles as shown in Fig. S1(a). Such a strong dependence was not seen for $\nu=1/3,2/3$ and $1/4$ [Fig. S1(b) (c) and (d), respectively], where the twist angle at which the crossover occurs shifts by only $\sim 0.5^{\circ}$.
Figure S1: Total ground-state energy for dielectric constant values $\varepsilon=5,10,15$ (first, second, and third columns from the left) in units of $e^{2}/(\varepsilon\ell_{B})$ for fillings (a) $3/4$, (b) $1/3$, and (c) $2/3$ (d) $1/4$.

### 10 Mar 2026 v2

@@ I.3 Effective potential
$\Pi_{+}=i\frac{\sqrt{2}\hbar}{\ell_{B}}\,a^{\dagger},\qquad\Pi_{-}=-i\frac{\sqrt{2}\hbar}{\ell_{B}}\,a.$ (S23)
@@
$\langle 1X^{\prime}|e^{i\bm{G}\cdot\bm{r}}|0X\rangle=i\frac{G_{-}\ell_{B}}{\sqrt{2}}\,\langle 0X^{\prime}|e^{i\bm{G}\cdot\bm{r}}|0X\rangle,$ (S24)
@@ II Landau-level operators
The following relations are useful and consistent with electrons in a magnetic field $\bm{B}=(0,0,-B)$ (equivalently, holes in $\bm{B}=(0,0,B)$) for $B>0$ [60]:
$\Pi_{x}$ $=p_{x}+eA_{x},\quad\Pi_{y}=p_{y}+eA_{y},\quad[\Pi_{i},\Pi_{j}]=+i\hbar^{2}\ell_{B}^{-2}\epsilon_{ij},\quad\ell_{B}=\sqrt{\hbar/eB},$ (S28)
$a$ $=i\frac{\ell_{B}}{\hbar\sqrt{2}}\left(\Pi_{x}+i\Pi_{y}\right),\quad a^{\dagger}=-i\frac{\ell_{B}}{\hbar\sqrt{2}}\left(\Pi_{x}-i\Pi_{y}\right),\quad[a,a^{\dagger}]=1,[a^{\dagger}a,a^{\dagger}]=a^{\dagger},$ (S29)
$\Pi_{x}$ $=i\frac{\hbar}{\ell_{B}\sqrt{2}}(a-a^{\dagger}),\quad\Pi_{y}=\frac{\hbar}{\ell_{B}\sqrt{2}}(a^{\dagger}+a),$ (S30)
$H_{0}$ $=\hbar\omega_{c}\left(a^{\dagger}a+\frac{1}{2}\right),\quad\omega_{c}=eB/m^{*},$ (S31)
$\bm{R}$ $=\bm{r}-\frac{\ell_{B}^{2}}{\hbar}\hat{z}\times\bm{\Pi},\quad[R_{i},R_{j}]=-i\ell_{B}^{2}\epsilon_{ij},\quad[R_{i},\Pi_{j}]=0,\quad[R_{i},H_{0}]=0,$ (S32)
$b$ $=\frac{1}{\sqrt{2}\ell_{B}}(R_{x}+iR_{y}),\quad b^{\dagger}=\frac{1}{\sqrt{2}\ell_{B}}(R_{x}-iR_{y}),\quad[b,b^{\dagger}]=1,\quad[b,a]=[b^{\dagger},a]=0,$ (S33)
$\bm{K}$ $=\bm{\Pi}+\hbar\ell_{B}^{-2}(\hat{z}\times\bm{r})=+\hbar\ell_{B}^{-2}(\hat{z}\times\bm{R}),\quad[K_{i},K_{j}]=-i\epsilon_{ij}\hbar^{2}\ell_{B}^{-2},$ (S34)
@@ III Commensurability of CDW and moiré patterns in a magnetic field
Let $n_{\phi}^{M}$ and $n_{\phi}^{\mathrm{CDW}}$ denote the magnetic flux per moiré and CDW unit cell, respectively. The topological non-trivial winding of the texture leads to $n_{\phi}^{M}=1$. We focus on commensurate periodicities that break translation symmetry while preserving $C_{6}$ rotational symmetry, implying a larger real-space CDW unit-cell area and $n_{\phi}^{\mathrm{CDW}}>n_{\phi}^{M}$. Assuming rational flux per CDW unit cell,
@@ IV Particle–hole transformation
$H_{\text{ad}}\rightarrow H_{\text{ad}}$ $=\frac{(\hat{\bm{p}}+e\bm{\mathcal{A}})^{2}}{2m^{*}}-V^{*}(\bm{r})$ (S45)
$=\frac{(\hat{\bm{p}}+e\bm{\mathcal{A}}_{0})^{2}}{2m^{*}}+\frac{1}{2m^{*}}\{(\hat{\bm{p}}+e\bm{\mathcal{A}}_{0}),\delta\bm{\mathcal{A}}(\bm{r})\}+\frac{(\delta\bm{\mathcal{A}}(\bm{r}))^{2}}{2m^{*}}-U_{\text{ad}}(\bm{r}).$ (S46)
@@ V Hartree–Fock matrix elements
We evaluate the exchange term analytically and implement the resulting closed-form expression in numerics [31] (Note: numerical implementations can also be found [33]). Total, single-particle, and interaction energies are
$E_{\mathrm{sp}}\equiv\mathrm{tr}(H_{0}\rho),\qquad E_{\mathrm{tot}}=\mathrm{tr}(H_{0}\rho)+\frac{1}{2}\,\mathrm{tr}(\Sigma^{\rm HF}\rho),\qquad E_{\mathrm{int}}=E_{\mathrm{tot}}-E_{\mathrm{sp}}.$ (S57)
@@ V.1 Density matrix and relations
where we used $\langle n^{\prime}X^{\prime}|e^{i\bm{q}\cdot\bm{r}}|nX\rangle=\delta_{X^{\prime},\,X+\ell_{B}^{2}q_{y}}\,e^{iq_{x}(X+X^{\prime})/2}\,F_{n^{\prime}n}(\bm{q})$ (see above) and the guiding-center density matrices
@@
In the CDW ground state characterized by reciprocal lattice vectors $\bm{G}$, we have
@@ V.2 Sum rule for the lowest Landau level
the usual LLL sum rule [31, 58, 57]. With Landau-level mixing, $\sum_{\bm{G}}|\Delta_{00}(\bm{G})|^{2}<\nu$.
@@ VI Lowest-Landau-level weight
To quantify how much a self-consistent solution resides in the LLL, define the lowest Landau-level weight $w_{\mathrm{LLL}}$:
@@ VII.1 Plane-wave matrix elements
$\langle k_{x}^{\prime}k_{y}^{\prime};n^{\prime}i^{\prime}|e^{i\bm{G}\cdot\bm{r}}|k_{x}k_{y};ni\rangle$ $=\delta_{k_{y},k_{y}^{\prime}}\,\delta_{k_{x},k_{x}^{\prime}}\,F_{n^{\prime}n}(\bm{G})\,e^{-iQ_{2}\,b\,k_{x}\ell_{B}^{2}}\,e^{i\ell_{B}^{2}G_{x}(k_{y}+G_{y}/2)}\,e^{iG_{x}\ell_{B}^{2}Q_{2}i}\,\delta_{\,i^{\prime},\,(i+b)\bmod j}.$ (S71)
@@
$\langle k_{x}^{\prime}k_{y}^{\prime};n^{\prime}i^{\prime}|e^{i\bm{G}\cdot\bm{r}}|k_{x}k_{y};ni\rangle$ $=\delta_{k_{y},k_{y}^{\prime}}\,\delta_{k_{x},k_{x}^{\prime}}\,\delta_{\,i^{\prime},\,(i+b)\bmod j}\,F_{n^{\prime}n}(\bm{G})\,e^{-2\pi i\,b\,\tilde{k}_{x}/j}\,e^{2\pi i\,\tilde{G}_{x}(\tilde{k}_{y}+b/2)/n_{\phi}}\,e^{2\pi i\,\tilde{G}_{x}i/n_{\phi}}.$ (S72)
@@
$S=\frac{L}{Q_{2}\,j\,\ell_{B}^{2}}=\frac{L\,Q_{0}\,p}{2\pi\,j\,q}.$ (S73)
@@ VII.3 Chern number
$|u_{\bm{k}};n,i\rangle=e^{-i\bm{k}\cdot\hat{\bm{r}}}|\bm{k};n,i\rangle\qquad\Longrightarrow\qquad\langle u_{\bm{k}};n,i|u_{\bm{k}^{\prime}};n^{\prime},i^{\prime}\rangle=\langle\bm{k};n,i|e^{i(\bm{k}-\bm{k}^{\prime})\cdot\bm{r}}|\bm{k}^{\prime};n^{\prime},i^{\prime}\rangle.$ (S74)
@@
$\langle\bm{k};n,i|e^{i(\bm{k}-\bm{k}^{\prime})\cdot\bm{r}}|\bm{k}^{\prime};n^{\prime},i^{\prime}\rangle$ $=\delta_{i,i^{\prime}}F_{nn^{\prime}}(\bm{k}-\bm{k}^{\prime})e^{-i\pi(\tilde{k}_{x}+\tilde{k}_{x}^{\prime})(\tilde{k}_{y}-\tilde{k}_{y}^{\prime})/j},\quad k_{x}=\tilde{k}_{x}Q_{0}p/(qj),\quad k_{y}=\tilde{k}_{y}Q_{2}.$ (S75)
@@
$|\bm{k},\alpha\rangle=\sum_{n,i}C^{\alpha}_{n,i}|\bm{k};n,i\rangle,$ $|u_{\bm{k}\alpha}\rangle=\sum_{n,i}C^{\alpha}_{n,i}e^{-i\bm{k}\cdot\hat{\bm{r}}}|\bm{k};n,i\rangle,$ $\Longrightarrow$ $\langle u_{\bm{k}\alpha}|u_{\bm{k}^{\prime}\alpha^{\prime}}\rangle=(\bm{C}^{\alpha}_{\bm{k}})^{\dagger}\mathrm{PW}(\bm{k},\bm{k}^{\prime})\bm{C}^{\alpha^{\prime}}_{\bm{k}^{\prime}},$ (S76)
@@
VIII Dielectric-constant-dependent phase transition at $\nu_{h}=3/4$
Here we discuss the dielectric-constant dependence of the ground-state energy at $\nu_{h}=3/4$. We find a pronounced $\varepsilon$-dependence of the energetically preferred lattice pinning configuration. For stronger Coulomb interactions (smaller $\varepsilon$), the $C_{6}$-symmetry-broken state in which the electron is pinned to the MX or XM site has lower energy for all twist angles in $2.5^{\circ},4.5^{\circ}]$, whereas for weaker interactions (larger $\varepsilon$) the transition shifts towards smaller twist angles as shown in Fig. S1(a). Such a strong dependence was not seen for $\nu_{h}=1/3$, $2/3$, and $1/4$ [Fig. S1(b), (c), and (d), respectively], where the twist angle at which the crossover occurs shifts by only $\sim 0.5^{\circ}$.
Figure S1: Total ground-state energy for dielectric constant values $\varepsilon=15,10,5$ (first, second, and third columns from the left) in units of $e^{2}/(4\pi\varepsilon\varepsilon_{0}\ell_{B})$ for fillings (a) $3/4$, (b) $1/3$, (c) $2/3$, and (d) $1/4$.